# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 5.1 Perfect production as extension of the market price
For $i=T-1$, (22) reduces to $0\leq C_{i+1}^{\prime,\theta}=\xi_{i+1}\cdot S_{i+1}+\tilde{Z}^{\xi}_{i+1}$ or equivalently $\phi^{\prime}_{i+1}\cdot S_{i+1}+\tilde{Z}^{\phi^{\prime}}_{i+1}\leq\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}$. As ϕ+superscriptitalic-$\phi^{\prime}\in\mathcal{R}^{+}$ (by Lemma 5.1) and $\theta\in\mathcal{R}^{\prime}$ are production strategies for $\mathcal{L}(\phi)$, they have the same cash flows for any $t\in D\setminus\{0,\dots,T\}$, so $\xi$ is self-financing. Thus, we can apply consistency of tradables (5) successively backward to get for any $i\leq t<i+1$ that $v_{t}(\phi^{\prime})\leq v_{t}(\theta)$, i.e. $v_{t}(\xi)\geq 0$. Thus, we can apply consistency with the tradables as in Definition 4.3 to conclude that $C_{i}\leq v_{i}(\xi)=v_{i}(\theta)-v_{i}(\phi^{\prime})$ and thus $\bar{v}_{i}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))=v_{i}(\theta)-C_{i}\geq v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Inserting this into (22) with $i+1$ replaced by $i$ gives $0\leq C_{i}^{\prime,\theta}=\xi_{i}\cdot S_{i}+\tilde{Z}^{\xi}_{i}$ for $i=T-1$. The proof proceeds recursively. ∎
@@
Let the almost sure fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out, and illiquid assets $\psi=0$. Assume consistency within the tradables in $\mathcal{R}$ (Definition 4.3). Then the following are equivalent:
@@
    For any strategy ϕ+italic-$\phi\in\mathcal{R}^{+}$ on $D[0,T]$ with zero incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}$, the value of the liability $\mathcal{L}(\phi)$ with $\bar{v}_{T}(\mathcal{L}(\phi))=0$ is $\bar{v}_{i}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$ for $i\in\{0,\ldots,T-1\}$.
@@
Consistency of the tradables with the financiability condition as in Definition 4.3 in principle allows investments in self-financing strategies of tradables to satisfy the financiability condition. If this is the case, the proof of Lemma 5.2 can be used to show that any production strategy $\theta\in\mathcal{R}^{\prime}$ with capital amounts $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ for the liability $\mathcal{L}(\phi)$ adjusted by potentially decreased capital amounts $\mathcal{C}^{*}$ has production cost equal to the market price, i.e. $\bar{v}_{i}^{\theta,0,\mathcal{C}^{*}}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Intuitively, capital is raised “uselessly” without increasing the production cost. It can be argued that this should be prevented by making the financiability condition stricter such that a higher return than the return of self-financing strategies is required, except of course for the investment $0\rightarrow 0$. This may also be plausible given the additional risks for investors when investing in an insurance company instead of in tradables. In this case, the definition of neutrality to the tradables from Definition 4.3 needs to be adjusted to at least exclude the capital investment $0\rightarrow 0$.
@@ 5.2 Adding short positions
We show in the following that the production cost is additive when adding a liability $\mathcal{L}(\phi)$ for ϕ+italic-$\phi\in\mathcal{R}^{+}$ as in Definition 3.3 with outgoing cash flows $X_{t}$ and zero incoming cash flows $\tilde{Z}_{t}=0$ to a liability $\mathcal{L}$ with terminal value $Y_{T}$. We assume the almost sure fulfillment condition and that portfolios with negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 3.3). Let $\theta\in\mathcal{R}^{\prime}$ with capital $\mathcal{C}=(C_{i})_{0}^{T-1}$ and illiquid assets $\psi$ be a production strategy for the liabilities $\mathcal{L}$. We show that $\theta+\phi^{\prime}$ for $\phi^{\prime}_{t}=1_{\{t\leq\tau\}}\cdot\phi_{t}$ with the same capital and illiquid assets is a production strategy for $\mathcal{L}+\mathcal{L}(\phi)$ with the same terminal value $Y_{T}$ and with production cost
@@ 5.3 Production with illiquid assets
Recall that we allow for a portfolio $\psi$ of “fully illiquid” assets in the production, which we assume are “held to maturity” (HTM). In particular, in contrast to a strategy of tradables, no additional illiquid assets can be bought, and only the incoming cash flows $\tilde{Z}_{t}^{\psi}$ of the illiquid assets at the given date $t$ can be used in the production and not the proceeds from selling part of the assets. Similarly, cash flows of the illiquid assets after date $T$ have an impact on the production only potentially when the assets are transferred at date $T$ to the capital investors, for whom their value is likely less than if the cash flows came from tradables. Holding assets to maturity thus potentially restricts the range of available production strategies, so a priori, the resulting value of the liabilities is at least as large as when trading assets is possible.
The treatment of the illiquid assets in the production is identical to how incoming cash flows (premiums) from the liabilities $\mathcal{L}$ are treated. For this reason, alternatively, we could view the object to be produced to be the liabilities $\mathcal{L}$ together with the illiquid assets $\psi$, effectively by adding the illiquid asset cash flows to the incoming premium cash flows. A similar type of “cash flow offsetting” that is common in practice is to value the insurance liabilities “net” of the cash flows of the outgoing reinsurance covers.
In the balance sheet value $A_{i}^{\prime}$ of the assets at a date $i$ according to Definition 4.1, no value is assigned to the future cash flows of the illiquid assets $\psi$, in line with the approach for production. At the same time, these future cash flows plausibly reduce the production cost on the liability side. By how much? One can conjecture that this reduction is equal to the value $v_{i}(\psi)$ of the outstanding cash flows in the case that $\psi$ is a suitable “static” investment strategy of tradables instead of illiquid assets, i.e. $\psi_{\gamma(t)}=\psi_{t}$ for any $t\in D$, outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, no incoming cash flows, and, for simplicity, terminal value $v_{T}(\psi)=0$. The following proposition shows this for financiability conditions that are neutral to the tradables and for general fulfillment conditions and strategies in $\mathcal{R}^{+}$. The proposition can alternatively be formulated for the almost sure fulfillment condition and strategies in $\mathcal{R}^{\prime}$ using suitable assumptions. Note that, for tradables of value $v_{i}(\psi)$ at date $i$, a different production strategy with lower production cost might be available.
@@
Let the financiability condition be neutral to the tradables (Definition 4.3). Given liabilities $\mathcal{L}$ on $D\equiv D[0,T]$ with terminal value $Y_{T}$, let ϕ+italic-$\phi\in\mathcal{R}^{+}$ with capital $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ be a production strategy for the liabilities $\mathcal{L}$ with no illiquid assets and production cost $\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})$. Let $\psi\in\mathcal{R}^{+}$ be an investment strategy of tradables with outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, no incoming cash flows, $\psi_{\gamma(t)}=\psi_{t}\geq 0$ on $D$, and $v_{T}(\psi)=0$. For $i\in\{0,\ldots,T-1\}$, let the investment strategy $\xi\equiv\xi^{(i)}\in\mathcal{R}^{+}$ of tradables be defined for $t\in D$ with $i<t<i+1$ such that $\xi_{\gamma(i)}=0$ and $\xi_{\gamma(t)}\cdot S_{t}=\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$ for $i<t<i+1$. Then, the strategy $\phi+\xi$ with capital $\mathcal{C}^{*}=(C_{i}+v_{i}(\psi))_{i=0}^{T-1}$ and “illiquid assets” $\psi$ is a production strategy for $\mathcal{L}$ with production cost $\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$.
@@
Let $i\in\{0,\ldots,T-1\}$. For $\phi+\xi\in\mathcal{R}^{+}$, condition (15) holds at $t\in D$ with $i<t<i+1$ because, using (15) for $\phi$, we have $(\phi+\xi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{(\phi+\xi)+\psi+\mathcal{L}}-X_{t}^{\mathcal{L}}=(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi+\mathcal{L}}-X_{t}^{\mathcal{L}})+(\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi})=(\phi+\xi)_{\gamma(t)}\cdot S_{t}$. We denote the value of the assets and liabilities at date $i+1$ by $A_{i+1}^{\prime}$ and $L_{i+1}$ for $\phi$, and by $A_{i+1}^{*\prime}$ and $L_{i+1}^{*}$ for $\phi+\xi$. Using that $\psi_{i+1}\cdot S_{i+1}=v_{i+1}(\psi)$ by assumption in the following last equality sign, we get $A_{i+1}^{*\prime}-L_{i+1}^{*}=(\phi+\xi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi+\xi+\psi+\mathcal{L}}-X_{i+1}^{\mathcal{L}}-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}+\bar{v}_{i+1}^{\phi,0,\mathcal{C}}(\mathcal{L})-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})-v_{i+1}(\psi)$. We proceed recursively backward in time and show that the last three terms together are equal to zero, i.e. $A_{i+1}^{*\prime}-L_{i+1}^{*}=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. For $i=T-1$, this holds because $\bar{v}_{T}^{\phi,0,\mathcal{C}}(\mathcal{L})=Y_{T}=\bar{v}_{T}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})$ and $v_{T}(\psi)=0$ by assumption. Then, because $(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}\geq 0$, the fulfillment condition holds for $\phi+\xi$ and for the capital payoff $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. The strategy $\xi+\psi\in\mathcal{R}^{+}$ is self-financing by construction: for $i<t<i+1$, we have $(\xi+\psi)_{\gamma(t)}\cdot S_{t}=(\xi+\psi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$. Neutrality of the financiability condition to the tradables then implies that $C_{i}^{*}=C_{i}+v_{i}(\xi+\psi)\rightarrow C_{i+1}^{*\prime}$ satisfies the fulfillment condition. Thus, for the production cost at date $i$, we get $\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=v_{i}(\phi+\xi)-C_{i}^{*}=v_{i}(\phi+\xi)-C_{i}-v_{i}(\xi+\psi)=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$ using $v_{i}(\xi)=0$ as $\xi_{\gamma(i)}=0$. Hence, we can proceed recursively. ∎
6 Valuation of insurance liabilities under Solvency II and the SST

### v2

@@ 4.1 Valuation in the framework as extension of market price valuation
For $i=T-1$, (22) reduces to $0\leq C_{i+1}^{\prime,\theta}=\xi_{i+1}\cdot S_{i+1}+\tilde{Z}^{\xi}_{i+1}$ or equivalently $\phi^{\prime}_{i+1}\cdot S_{i+1}+\tilde{Z}^{\phi^{\prime}}_{i+1}\leq\theta_{i+1}\cdot S_{i+1}+\tilde{Z}^{\theta}_{i+1}$. As $\phi^{\prime}\in\mathcal{R}^{\geq 0}$ (by Lemma 4.1) and $\theta\in\mathcal{R}^{\prime}$ are production strategies for $\mathcal{L}(\phi)$, they have the same cash flows for any $t\in D\setminus\{0,\dots,T\}$, so $\xi$ is self-financing. Thus, we can apply consistency of tradables (5) successively backward to get for any $i\leq t<i+1$ that $v_{t}(\phi^{\prime})\leq v_{t}(\theta)$, i.e. $v_{t}(\xi)\geq 0$. Thus, we can apply consistency with the tradables as in Definition 3.6 to conclude that $C_{i}\leq v_{i}(\xi)=v_{i}(\theta)-v_{i}(\phi^{\prime})$ and thus $\bar{v}_{i}^{\theta,0,\mathcal{C}}(\mathcal{L}(\phi))=v_{i}(\theta)-C_{i}\geq v_{i}(\phi^{\prime})=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Inserting this into (22) with $i+1$ replaced by $i$ gives $0\leq C_{i}^{\prime,\theta}=\xi_{i}\cdot S_{i}+\tilde{Z}^{\xi}_{i}$ for $i=T-1$. The proof proceeds recursively. ∎
@@
Let the full fulfillment condition apply, portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out, and illiquid assets $\psi=0$. Assume consistency within the tradables in $\mathcal{R}$ (Definition 3.6). Then the following are equivalent:
@@
    For any strategy $\phi\in\mathcal{R}^{\geq 0}$ on $D[0,T]$ with zero cash inflows $\tilde{Z}_{t}=0$ and cash outflows $X_{t}$, the value of the liability $\mathcal{L}(\phi)$ with $\bar{v}_{T}(\mathcal{L}(\phi))=0$ is $\bar{v}_{i}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$ for $i\in\{0,\ldots,T-1\}$.
@@
Requiring a financiability condition to be consistent with the tradables as in Definition 3.6 does not preclude that investments in self-financing strategies of tradables satisfy the financiability condition. If self-financing strategies of tradables satisfy the financiability condition, the proof of Lemma 4.2 can be used to show that any production strategy $\theta\in\mathcal{R}^{\prime}$ with capital amounts $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ for the liability $\mathcal{L}(\phi)$, potentially adjusted by replacing $\mathcal{C}$ with lower capital amounts $\mathcal{C}^{*}$, has “minimal” production cost equal to the market price, i.e. $\bar{v}_{i}^{\theta,0,\mathcal{C}^{*}}(\mathcal{L}(\phi))=v_{i}(\phi)\cdot 1_{\{i<\tau\}}$. Intuitively speaking, this means in particular that capital can be raised, even when it would not be required, without increasing the production cost. It can be argued that this should be prevented by making the financiability condition stricter. That is, such that capital investments require a higher return than the return from self-financing strategies (except of course for the investment $0\rightarrow 0$). This may also be plausible given the additional risks investors may face from investing in an insurance company instead of in tradables. In this case, the financiability condition cannot also be neutral to the tradables according to Definition 3.6, at least not for $C_{i}=0$ and $C_{i+1}^{\prime}=0$.
Theorem 4.3 is limited to perfect fulfillment. With a weaker fulfillment condition, we can for example not in general expect that the liability $\mathcal{L}(\phi)=0$ cannot be produced with a production strategy with negative production cost. For perfect fulfillment, on the other hand, further normative properties can be expected to hold. In particular, the following corollary shows that a liability with “non-negative cash flows” has a non-negative value.
Corollary 4.5.
Let the full fulfillment condition apply and portfolios containing negative units in $\mathcal{R}^{\prime}$ be available for production with close out. Assume consistency within the tradables in $\mathcal{R}$ and with the financiability condition. Let the liability $\mathcal{L}$ with terminal value $Y_{T}=0$ and the illiquid assets $\psi$ be such that “cash flows are non-negative” in the sense that, for all $t\in D\equiv D[0,T]$,
$X_{t}^{\mathcal{L}}-\tilde{Z}_{t}^{\psi+\mathcal{L}}\geq 0$ (23)
Then the liability $\mathcal{L}$ has non-negative value, $\bar{v}_{i}^{\psi}(\mathcal{L})\geq 0$ for $i\in\{0,\ldots,T-1\}$.
Proof.
Given an arbitrary production strategy ϕitalic-$\phi\in\mathcal{R}^{\prime}$ for $\mathcal{L}$, we construct a production strategy $\theta=\phi+\xi\in\mathcal{R}^{\prime}$ with $\xi\in\mathcal{R}^{\geq 0}$, capital $\mathcal{C}(\theta)$, and no illiquid assets for the liability $\mathcal{L}(\phi)=0$ with terminal value 00 such that $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq\bar{v}_{i}^{\theta,0,\mathcal{C}(\theta)}(0)$ for $i\in\{0,\ldots,T-1\}$. It then follows from the definition of the value and Theorem 4.3 that $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq\bar{v}_{i}(0)=0$, so the claim follows from the definition of the value $\bar{v}_{i}^{\psi}(\mathcal{L})$.
We construct the production strategy $\theta$ recursively backward in time for $i\in\{0,\ldots,T-1\}$, forward in time within each one-year period from date $i$ to $i+1$, such that $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq\bar{v}_{i}^{\theta,0,\mathcal{C}(\theta)}(0)$. At date $i+1=T$, the inequality holds as $\bar{v}_{T}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=0=\bar{v}_{T}^{\theta,0,\mathcal{C}(\theta)}(0)$ by assumption. From $i$ to $i+1$, we define $\xi_{\gamma(i)}=0$ and, at $\gamma(i)\leq t<i+1$, given $\xi_{t}$, we select $\xi_{\gamma(t)}\in\mathcal{R}^{\geq 0}$ such that
$\xi_{\gamma(t)}\cdot S_{t}=\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi}+X_{t}^{\mathcal{L}}-\tilde{Z}_{t}^{\psi+\mathcal{L}}\geq 0$ (24)
which is possible as $X_{t}^{\mathcal{L}}-\tilde{Z}_{t}^{\psi+\mathcal{L}}\geq 0$ by assumption. Then, $\theta_{\gamma(t)}\cdot S_{t}=\phi_{\gamma(t)}\cdot S_{t}+\xi_{\gamma(t)}\cdot S_{t}$ is non-negative as $\phi_{\gamma(t)}\cdot S_{t}\geq 0$ and $\xi_{\gamma(t)}\cdot S_{t}\geq 0$ and, using (a) from Definition 3.7 for $\phi$, equal to
$\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi+\psi+\mathcal{L}}-X_{t}^{\mathcal{L}}+\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi}+X_{t}^{\mathcal{L}}-\tilde{Z}_{t}^{\psi+\mathcal{L}}=\theta_{t}\cdot S_{t}+\tilde{Z}_{t}^{\theta}-0$ (25)
so that (a) from Definition 3.7 holds for $\theta$. At date $i+1$, we denote the assets and liabilities according to Definition 3.1 by $A_{i+1}^{\prime,\phi}$ and $L_{i+1}^{\phi}$ for $\mathcal{L}$ and the strategy $\phi$, and as $A_{i+1}^{\prime,\theta}$ and $L_{i+1}^{\theta}$ for the liability 00 and the strategy $\theta$. Using the respective definitions, we can write $A_{i+1}^{\prime,\theta}-L_{i+1}^{\theta}$ as
$A_{i+1}^{\prime,\phi}-L_{i+1}^{\phi}+(\xi_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi})+(X_{i+1}^{\mathcal{L}}-\tilde{Z}_{i+1}^{\psi+\mathcal{L}})+(\bar{v}_{i+1}^{\phi,\psi,\mathcal{C}}(\mathcal{L})-\bar{v}_{i+1}^{\theta,0,\tilde{\mathcal{C}}}(0))$ (26)
By the various properties and assumptions, $A_{i+1}^{\prime,\theta}-L_{i+1}^{\theta}\geq A_{i+1}^{\prime,\phi}-L_{i+1}^{\phi}\geq 0$, so the full fulfillment condition is satisfied for $\theta$. Hence, also, for the return to the capital provider, $C_{i+1}^{\prime,\theta}\geq C_{i+1}^{\prime,\phi}$ with the analogue notation, hence by the properties of the financiability condition, $C_{i}^{\theta}\geq C_{i}^{\phi}\equiv C_{i}$. We thus get for the production cost $\bar{v}_{i}^{\theta,0,\mathcal{C}(\theta)}(0)$ date $i$, using that $v_{i}(\theta)=\theta_{\gamma(i)}\cdot S_{i}=\phi_{\gamma(i)}\cdot S_{i}=v_{i}(\phi)$,
$\bar{v}_{i}^{\theta,0,\mathcal{C}(\theta)}(0)=v_{i}(\theta)-C_{i}^{\theta}\leq v_{i}(\phi)-C_{i}=\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ (27)
The proof then proceeds recursively. ∎
@@ 4.2 Adding short positions
We show in the following that the production cost is additive when adding a liability $\mathcal{L}(\phi)$ for $\phi\in\mathcal{R}^{\geq 0}$ as in Definition 2.4 with cash outflows $X_{t}$ and zero cash inflows $\tilde{Z}_{t}=0$ to a liability $\mathcal{L}$ with terminal value $Y_{T}$. We assume the full fulfillment condition and that portfolios with negative units in $\mathcal{R}^{\prime}$ are available for production with close out (Definition 2.4). Let $\theta\in\mathcal{R}^{\prime}$ with capital $\mathcal{C}=(C_{i})_{0}^{T-1}$ and illiquid assets $\psi$ be a production strategy for the liabilities $\mathcal{L}$. We show that $\theta+\phi^{\prime}$ for $\phi^{\prime}_{t}=1_{\{t\leq\tau\}}\cdot\phi_{t}$ with the same capital and illiquid assets is a production strategy for $\mathcal{L}+\mathcal{L}(\phi)$ with the same terminal value $Y_{T}$ and with production cost
@@ 4.3 Production with illiquid assets
Recall that we allow for a portfolio $\psi$ of “fully illiquid” assets in the production, which we assume are “held to maturity” (HTM). In particular, in contrast to a strategy of tradables, no additional illiquid assets can be bought, and only the cash inflows $\tilde{Z}_{t}^{\psi}$ of the illiquid assets at the given date $t$ can be used in the production and not the proceeds from selling part of the assets. Similarly, cash flows of the illiquid assets after date $T$ have an impact on the production only potentially when the assets are transferred at date $T$ to the capital investors, for whom their value is likely less than if the cash flows came from tradables. Holding assets to maturity thus potentially restricts the range of available production strategies, so a priori, the resulting value of the liabilities is at least as large as when trading assets is possible.
The treatment of the illiquid assets in the production is identical to how cash inflows (premiums) from the liabilities $\mathcal{L}$ are treated. For this reason, alternatively, we could view the object to be produced to be the liabilities $\mathcal{L}$ together with the illiquid assets $\psi$, effectively by adding the illiquid asset cash flows to the premium cash inflows. A similar type of “cash flow offsetting” that is common in practice is to value the insurance liabilities “net” of the cash flows of the outgoing reinsurance covers.
In the balance sheet value $A_{i}^{\prime}$ of the assets at a date $i$ according to Definition 3.1, no value is assigned to the future cash flows of the illiquid assets $\psi$, in line with the approach for production. At the same time, these future cash flows plausibly reduce the production cost on the liability side. By how much? One can conjecture that this reduction is equal to the value $v_{i}(\psi)$ of the outstanding cash flows in the case that $\psi$ is a suitable “static” investment strategy of tradables instead of illiquid assets, i.e. $\psi_{\gamma(t)}=\psi_{t}$ for any $t\in D$, cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, no cash inflows, and, for simplicity, terminal value $v_{T}(\psi)=0$. The following proposition shows this for financiability conditions that are neutral to the tradables and for general fulfillment conditions and strategies in $\mathcal{R}^{\geq 0}$. The proposition can alternatively be formulated for the full fulfillment condition and strategies in $\mathcal{R}^{\prime}$ using suitable assumptions. Note that, for tradables of value $v_{i}(\psi)$ at date $i$, a different production strategy with lower production cost might be available.
@@
Let the financiability condition be neutral to the tradables (Definition 3.6). Given liabilities $\mathcal{L}$ on $D\equiv D[0,T]$ with terminal value $Y_{T}$, let $\phi\in\mathcal{R}^{\geq 0}$ with capital $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ be a production strategy for the liabilities $\mathcal{L}$ with no illiquid assets and production cost $\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})$. Let $\psi\in\mathcal{R}^{\geq 0}$ be an investment strategy of tradables with cash outflows $X_{t}=\tilde{Z}_{t}^{\psi}$, no cash inflows, $\psi_{\gamma(t)}=\psi_{t}\geq 0$ on $D$, and $v_{T}(\psi)=0$. For $i\in\{0,\ldots,T-1\}$, let the investment strategy $\xi\equiv\xi^{(i)}\in\mathcal{R}^{\geq 0}$ of tradables be defined for $t\in D$ with $i<t<i+1$ such that $\xi_{\gamma(i)}=0$ and $\xi_{\gamma(t)}\cdot S_{t}=\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$ for $i<t<i+1$. Then, the strategy $\phi+\xi$ with capital $\mathcal{C}^{*}=(C_{i}+v_{i}(\psi))_{i=0}^{T-1}$ and “illiquid assets” $\psi$ is a production strategy for $\mathcal{L}$ with production cost
$\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$ (29)
@@
Let $i\in\{0,\ldots,T-1\}$. For $\phi+\xi\in\mathcal{R}^{\geq 0}$, condition (17) holds at $t\in D$ with $i<t<i+1$ because, using (17) for $\phi$, we have $(\phi+\xi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{(\phi+\xi)+\psi+\mathcal{L}}-X_{t}^{\mathcal{L}}=(\phi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\phi+\mathcal{L}}-X_{t}^{\mathcal{L}})+(\xi_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi})=(\phi+\xi)_{\gamma(t)}\cdot S_{t}$. We denote the value of the assets and liabilities at date $i+1$ by $A_{i+1}^{\prime}$ and $L_{i+1}$ for $\phi$, and by $A_{i+1}^{*\prime}$ and $L_{i+1}^{*}$ for $\phi+\xi$. Using that $\psi_{i+1}\cdot S_{i+1}=v_{i+1}(\psi)$ by assumption in the following last equality sign, we get $A_{i+1}^{*\prime}-L_{i+1}^{*}=(\phi+\xi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\phi+\xi+\psi+\mathcal{L}}-X_{i+1}^{\mathcal{L}}-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}+\bar{v}_{i+1}^{\phi,0,\mathcal{C}}(\mathcal{L})-\bar{v}_{i+1}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})-v_{i+1}(\psi)$. We proceed recursively backward in time and show that the last three terms together are equal to zero, i.e. $A_{i+1}^{*\prime}-L_{i+1}^{*}=A_{i+1}^{\prime}-L_{i+1}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. For $i=T-1$, this holds because $\bar{v}_{T}^{\phi,0,\mathcal{C}}(\mathcal{L})=Y_{T}=\bar{v}_{T}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})$ and $v_{T}(\psi)=0$ by assumption. Then, because $(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}\geq 0$, the fulfillment condition holds for $\phi+\xi$ and for the capital payoff $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}+(\xi+\psi)_{i+1}\cdot S_{i+1}+\tilde{Z}_{i+1}^{\xi+\psi}$. The strategy $\xi+\psi\in\mathcal{R}^{\geq 0}$ is self-financing by construction: for $i<t<i+1$, we have $(\xi+\psi)_{\gamma(t)}\cdot S_{t}=(\xi+\psi)_{t}\cdot S_{t}+\tilde{Z}_{t}^{\xi+\psi}$. Neutrality of the financiability condition to the tradables then implies that $C_{i}^{*}=C_{i}+v_{i}(\xi+\psi)\rightarrow C_{i+1}^{*\prime}$ satisfies the fulfillment condition. Thus, for the production cost at date $i$, we get $\bar{v}_{i}^{\phi+\xi,\psi,\mathcal{C}^{*}}(\mathcal{L})=v_{i}(\phi+\xi)-C_{i}^{*}=v_{i}(\phi+\xi)-C_{i}-v_{i}(\xi+\psi)=\bar{v}_{i}^{\phi,0,\mathcal{C}}(\mathcal{L})-v_{i}(\psi)$ using $v_{i}(\xi)=0$ as $\xi_{\gamma(i)}=0$. Hence, we can proceed recursively. ∎
5 Valuation of insurance liabilities under Solvency II and in the SST